Tickets for a school play sell for $8 for floor seats & $6 for balcony seats. 72 tickets are sold, bringing in $516. How many of each ticket are sold? Must show how. Thank you
Answer by gokhalesujata(4)

Let there be x floor seats tickets sold and y balcony seats tickets sold
x+y=72 as 72 tickets were sold
Also 8x+6y=516
solving we get
x=72-y
substituting for x we get
8*(72-y)+6*y=516
576-8y+6y=516
576-516-2y=0
60=2y
y=30
x=72-30=42